Grade 12Algebra

Decomposition of H2O2 follows a first order reaction. In fifty minutes the concentration of H2O2 decreases from 0.5 to 0.125 M in one such decomposition. When the concentration of H2O2 reaches 0.05 M, the rate of formation of O2 will be

Narendra
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
As we know
K = (1/t) ln (ao/at)
K = (1/50) ln(4)
Rate of disappearance of H2O2 = rate of appearance of O2/(1/2)
r O2 = r H2O2 /2
 = 1/2 * 1/50 * ln 4 * 0.05
 = 6.8 * 10^-4